OpenStudy (anonymous):

(1/2)log[(3)(c^2-d^2)-log(3)(c-d)] write as single logarithm

OpenStudy (anonymous):

man you never quit do you?

OpenStudy (anonymous):

forget the base, it is unimportant so it just says "write as a single log"

OpenStudy (anonymous):

sry to bother, you don't need to help me if you don't want to

OpenStudy (anonymous):

first use \[\log(A)-\log(B)=\log(\frac{A}{B})\]

OpenStudy (anonymous):

i am just teasing. late at night and back by noon!

OpenStudy (anonymous):

yeah, I just want this to be over and I feel like an old dog now that can't learn new tricks

OpenStudy (anonymous):

so you get \[\log(\frac{c^2-d^2}{c-d})\]

OpenStudy (anonymous):

then factor and cancel on the inside of the log to get \[\frac{c^2-d^2}{c-d}=\frac{(c+d)(c-d)}{c-d}=c+d\]

OpenStudy (anonymous):

giving you \[\frac{1}{2}\log(c+d)\]

OpenStudy (anonymous):

can I ask, are you a math teacher? or you're just always here because you are good and it's enjoyable for you?

OpenStudy (anonymous):

now use that \[n\log(x)=\log(x^n)\] to get \[\frac{1}{2}\log(c+d)=\log((c+d)^{\frac{1}{2}})=\log(\sqrt{c+d})\]\]

OpenStudy (anonymous):

as for your question, i hate to see people have to struggle through questions like this with essentially no content and no meaning, and especially no understanding. it is just a bunch of ruled combined with some algebra rules http://www.unc.edu/depts/jomc/academics/dri/idog.html

OpenStudy (anonymous):

that's awesome. I want to spend time helping people on the writing part of this site, but there aren't nearly as many needing help!